Solve the following formula for y.
x=9y-5
I'm honestly not sure where to even start?

Answer by Alan3354(69443)

x=9y-5
Solve for y means get y on one side on the equal sign by itself.
x=9y-5
Add 5
x + 5 = 9y
9y = x + 5
Divide by 9
y = (x + 5)/9